Jasmine is cycling at 12 m s−1 when her bag falls off the back of her bicycle.The bag slides a distance of 9 m before coming to rest. Calculate the coefficient of friction between the bag and the road.


1
Expert's answer
2021-11-23T13:00:04-0500

The work-energy theorem states that the work done on an object by the net force is equal to the change in its kinetic energy:


"\dfrac{mv_2^2}{2}-\dfrac{mv_1^2}{2}=-F_{fr}\cdot s""F_{fr}=\mu m g"

"\dfrac{mv_2^2}{2}-\dfrac{mv_1^2}{2}=-\mu mg\cdot s"

"\mu=\dfrac{v_1^2-v_2^2}{2gs}"

Given "v_1=12\ m/s, v_2=0\ m/s, s=9m, g=9.81\ m/s^2."


"\mu=\dfrac{(12\ m/s)^2-(0\ m/s)^2}{2(9.81\ m/s^2)(9\ m)}=0.8155"

The coefficient of friction between the bag and the road is 0.8155.